# Discrete Probabilistic Inference as Control in Multi-path Environments

### v1

@@ 5.1 Probabilistic inference over discrete factor graphs
The probabilistic inference task in Buesing et al. (2020) consists in sequentially sampling the values of $d$ discrete random variables in a factor graph one at a time, with a fixed order. This makes the underlying MDP having a tree structure, eliminating the need for reward correction, as described in Section 2.2. We adapted this environment to have multiple trajectories leading to each terminating state by allowing sampling the random variables in any order. Details about the energy function are available in Section C.1.
In Figure 3, we show the performance of the different MaxEnt RL and GFlowNet algorithms on 4 different factor graph structures, as proposed by Buesing et al. (2020), with $d=6$ variables and where each variable can take one of 5 possible values. We observe that TB & PCL perform similarly, validating Proposition 3.2, and overall outperform all other methods. Similarly, we can see that DB & SQL also perform similarly as expected by Corollary B.1. Finally, although SAC is generally viewed as a strong algorithm for continuous control (Haarnoja et al., 2018b), we did not observe any significant improvement over DB/SQL.
@@ 5.2 Structure learning of Bayesian Networks
We also evaluated all algorithms on the task of learning the structure of Bayesian networks, using a Bayesian perspective (Deleu et al., 2022). Given a dataset of observations ${\mathcal{D}}$ from a joint distribution over $d$ random variables $X_{1},\ldots,X_{d}$, our objective is to approximate the posterior distribution $P(G\mid{\mathcal{D}})\propto P({\mathcal{D}}\mid G)P(G)$ over the DAG structures $G$ encoding the conditional independencies of $P(X_{1},\ldots,X_{d})$. The soft MDP is constructed as in (Deleu et al., 2022), where a DAG $G$ is constructed by adding one edge at a time, starting from a completely empty graph, while enforcing the acyclicity of the graph at each step of generation (i.e., an edge cannot be added if it would introduce a cycle).
Figure 4: Comparison of MaxEnt RL and GFlowNet algorithms on the Bayesian structure learning task, in terms of the Jensen-Shannon divergence between the terminating state distribution and the target posterior during training. Both experiments differ in the way the marginal likelihood $P({\mathcal{D}}\mid G)$ is computed, (left) using the BGe score (Geiger and Heckerman, 1994), (right) is the linear Gaussian score (Nishikawa-Toomey et al., 2023). Each curve represents the average JSD with 95% confidence interval over 20 random seeds.
Following Malkin et al. (2023), we consider here a relatively small task where $d=5$, so that the target distribution $P(G\mid{\mathcal{D}})$ can be evaluated analytically in order to compare it to our approximations given by MaxEnt RL and/or GFlowNets. Unlike in Section 5.1, we also included the modified DB loss and $\pi$-SQL in our comparison since all the states are valid DAGs. We observe in Figure 4 that again TB & PCL on the one hand, but also modified DB & $\pi$-SQL on the other hand perform very similarly to one another, empirically validating our equivalences established above. Despite a light search over hyperparameters, we found that SAC performs on average significantly worse than other methods, mainly due to instability during training.
@@ 5.3 Phylogenetic tree generation
Finally, we also compared these methods on the larger-scale task of parsimonious phylogenetic tree generation introduced by Zhou et al. (2024). Based on biological sequences of different species, the objective is to find phylogenetic trees over those species that require few mutations. A state of the soft MDP corresponds to a collection of trees over a partition of the species, and actions correspond to merging two trees together by adding a root node. Note that all the trees sampled this way have the same size, although they may have different number of trajectories leading to each of them (unlike Figure 3). The energy of a tree $T$ corresponds to the total number of mutations captured in $T$.
In Figure 5, we compare the performance of all methods on 6 datasets introduced by Zhou et al. (2024), in terms of the correlation between the terminating state log-probabilities $\log\pi(T)$ associated with the learned policy, and the (uncorrected) return $-{\mathcal{E}}(T)$, since we should ideally have
@@
Figure 5: Comparison of MaxEnt RL and GFlowNet algorithms on the phylogenetic tree generation task. (Left) Comparison of the performance in terms of the Pearson correlation between the terminating state log-probability and the return on 1000 randomly sampled trees. (Center) Correlation between the terminating state log-probability found with DB and the return, each point representing a tree, with a best linear fit line and its slope. (Right) Similarly for SQL. The correlation plots for all methods and all datasets are available in Section C.3.
@@ 6 Discussion
Similar to (Mohammadpour et al., 2024), we note that our results here are limited to the case where the soft MDP is deterministic, to match the standard assumptions made in the GFlowNet literature. Although some works have attempted to generalize GFlowNets to stochastic environments (Bengio et al., 2023; Pan et al., 2023b), there is no apparent consensus on how to guarantee the existence of an optimal policy $P_{F}$ whose terminating state distribution matches Equation 1 when an object may be generated in multiple ways (i.e., ${\mathcal{G}}$ is a DAG). Jiralerspong et al. (2024) introduced an extension to stochastic environments called EFlowNet where an optimal policy is guaranteed to exist, albeit limited to the case where the soft MDP has a tree structure and would therefore bypass the need for any reward correction. The assumption of determinism is generally not limiting though since the structure of the soft MDP is typically designed by an expert based on the problem at hand (depending on the distribution to be approximated).
@@
We saw in Section 2.2 that without correcting the reward function, the optimal policy $\pi_{\mathrm{MaxEnt}}^{*}$ has a terminating state distribution biased towards states with more complete trajectories leading to them. However, there are some situations where all the states have an equal number of trajectories leading to them. This is the case of the discrete factor graphs environments studied in Section 5.1 for example, where all terminating states can be accessed with exactly $n!$ trajectories since each of the $n$ variables can be assigned a value in any order. In this situation, one can apply MaxEnt RL with the original reward in Equation 2, without any correction, and still obtain a terminating state distribution equal to the Gibbs distribution, since this constant can be absorbed into the partition function. Just like the case where ${\mathcal{G}}$ is a tree, this must be considered as a very special case though, and it is generally recommended to always correct the reward with $P_{B}$.
@@
In all applications of GFlowNets involving the SubTB objective (Madan et al., 2023; Malkin et al., 2022) found in the literature, the forward transition probabilities $P_{F}^{\theta}$ and the state flow function $F_{\phi}$ have always been parametrized with separate models (with possibly a shared backbone) as in Section 3.2. However, thanks to the equivalence with PCL established in Proposition 3.2, it is actually possible to parametrize both functions using a single “Q-function” thanks to the Unified PCL perspective (Nachum et al., 2017); see also Section B.2 in the context of DB & SQL. A setting with separate networks is closer to the use of a dueling architecture as highlighted by Tiapkin et al. (2024). In our experiments, we observed that using DB with two separate networks performs overall similarly to SQL with a single Q-network, even though the latter requires fewer parameters to learn. A complete study of the effectiveness of this strategy compared to separate networks is left as future work.
@@
While this paper focused on discrete distributions, to match the well-studied setting in the GFlowNet literature, the reward correction in Theorem 3.1 may be extended to cases where ${\mathcal{X}}$, along with the state and action spaces of the soft MDP, are continuous spaces as conjectured by Tiapkin et al. (2024). We could establish similar connections with MaxEnt RL objectives by leveraging the extensions of GFlowNets to continuous spaces (Li et al., 2023; Lahlou et al., 2023). Interestingly, continuous GFlowNets have strong connections with diffusion models (Zhang et al., 2022a; Sendera et al., 2024), for which RL has been shown to be an effective training method (Fan and Lee, 2023; Black et al., 2023).
@@ Acknowledgements
We would like to thank Valentin Thomas, Michal Valko, Pierre Ménard, Daniil Tiapkin, and Sobhan Mohammadpour for helpful discussions and comments about this paper. This research was enabled in part by compute resources and software provided by Mila (mila.quebec).
Appendix
@@ Appendix A Reward correction
Recall from Ziebart [2010], Haarnoja et al. [2017] that the optimal policy maximizing Equation 3 is
@@ Appendix B Equivalences between MaxEnt RL & GFlowNet objectives
${}^{\star}$ [Haarnoja et al., 2017] FL-DB [Pan et al., 2023a] Proposition B.3 App. B.4
$\pi$${}^{\star}$ (Section 3.3) M-DB [Deleu et al., 2022] Proposition 3.3 App. B.3
@@ C.2 Structure learning of Bayesian Networks
A state of the soft MDP corresponds to a DAG $G$ over $d$ nodes, and the initial state is the empty graph over $d$ nodes. An action consists in adding a directed edge between two nodes, such that it is not already present in the graph, and it doesn’t introduce a cycle, garanteeing that all the states of the MDP are valid acyclic graphs; there is a special action indicating whether we want to terminate and transition to $s_{f}$.

### v2

@@ 5.1 Probabilistic inference over discrete factor graphs
The probabilistic inference task in Buesing et al. [2020] consists in sequentially sampling the values of $d$ discrete random variables in a factor graph one at a time, with a fixed order. This makes the underlying MDP having a tree structure, eliminating the need for reward correction, as described in Section 2.2. We adapted this environment to have multiple trajectories leading to each terminating state by allowing sampling the random variables in any order. Details about the energy function are available in Section C.1.
In Figure 3, we show the performance of the different MaxEnt RL and GFlowNet algorithms on 4 different factor graph structures, as proposed by Buesing et al. [2020], with $d=6$ variables and where each variable can take one of 5 possible values. We observe that TB & PCL perform similarly, validating Proposition 3.2, and overall outperform all other methods. Similarly, we can see that DB & SQL also perform similarly as expected by Corollary B.1. Finally, although SAC is generally viewed as a strong algorithm for continuous control [Haarnoja et al., 2018b], we did not observe any significant improvement over DB/SQL.
Figure 4: Comparison of MaxEnt RL and GFlowNet algorithms on the Bayesian structure learning task, in terms of the Jensen-Shannon divergence between the terminating state distribution and the target posterior during training. Both experiments differ in the way the marginal likelihood $P({\mathcal{D}}\mid G)$ is computed, (left) using the BGe score [Geiger and Heckerman, 1994], (right) is the linear Gaussian score [Nishikawa-Toomey et al., 2023]. Each curve represents the average JSD with 95% confidence interval over 20 random seeds.
@@ 5.2 Structure learning of Bayesian Networks
We also evaluated all algorithms on the task of learning the structure of Bayesian networks, using a Bayesian perspective [Deleu et al., 2022]. Given a dataset of observations ${\mathcal{D}}$ from a joint distribution over $d$ random variables $X_{1},\ldots,X_{d}$, our objective is to approximate the posterior distribution $P(G\mid{\mathcal{D}})\propto P({\mathcal{D}}\mid G)P(G)$ over the DAG structures $G$ encoding the conditional independencies of $P(X_{1},\ldots,X_{d})$. The soft MDP is constructed as in [Deleu et al., 2022], where a DAG $G$ is constructed by adding one edge at a time, starting from a completely empty graph, while enforcing the acyclicity of the graph at each step of generation (i.e., an edge cannot be added if it would introduce a cycle).
Following Malkin et al. [2023], we consider here a relatively small task where $d=5$, so that the target distribution $P(G\mid{\mathcal{D}})$ can be evaluated analytically in order to compare it to our approximations given by MaxEnt RL and/or GFlowNets. Unlike in Section 5.1, we also included the modified DB loss and $\pi$-SQL in our comparison since all the states are valid DAGs. We observe in Figure 4 that again TB & PCL on the one hand, but also modified DB & $\pi$-SQL on the other hand perform very similarly to one another, empirically validating our equivalences established above. Despite a light search over hyperparameters, we found that SAC performs on average significantly worse than other methods, mainly due to instability during training.
@@ 5.3 Phylogenetic tree generation
Finally, we also compared these methods on the larger-scale task of parsimonious phylogenetic tree generation introduced by Zhou et al. [2024]. Based on biological sequences of different species, the objective is to find phylogenetic trees over those species that require few mutations. A state of the soft MDP corresponds to a collection of trees over a partition of the species, and actions correspond to merging two trees together by adding a root node. Note that all the trees sampled this way have the same size, although they may have different number of trajectories leading to each of them (unlike Figure 3). The energy of a tree $T$ corresponds to the total number of mutations captured in $T$.
Figure 5: Comparison of MaxEnt RL and GFlowNet algorithms on the phylogenetic tree generation task. (Left) Comparison of the performance in terms of the Pearson correlation coefficient between the terminating state log-probability and the return on 1000 randomly sampled trees. (Center) Correlation between the terminating state log-probability found with DB and the return, each point representing a tree, with a best linear fit line and its slope. (Right) Similarly for SQL. The correlation plots for all methods and all datasets are available in Section C.3.
In Figure 5, we compare the performance of all methods on 6 datasets introduced by Zhou et al. [2024], in terms of the correlation between the terminating state log-probabilities $\log\pi(T)$ associated with the learned policy, and the (uncorrected) return $-{\mathcal{E}}(T)$, since we should ideally have
@@ 6 Discussion
Similar to [Mohammadpour et al., 2024], we note that our results here are limited to the case where the soft MDP is deterministic, to match the standard assumptions made in the GFlowNet literature. Although some works have attempted to generalize GFlowNets to stochastic environments [Bengio et al., 2023, Pan et al., 2023b], there is no apparent consensus on how to guarantee the existence of an optimal policy $P_{F}$ whose terminating state distribution matches Equation 1 when an object may be generated in multiple ways (i.e., ${\mathcal{G}}$ is a DAG). Jiralerspong et al. [2024] introduced an extension to stochastic environments called EFlowNet where an optimal policy is guaranteed to exist, albeit limited to the case where the soft MDP has a tree structure and would therefore bypass the need for any reward correction. The assumption of determinism is generally not limiting though since the structure of the soft MDP is typically designed by an expert based on the problem at hand (depending on the distribution to be approximated).
@@
We saw in Section 2.2 that without correcting the reward function, the optimal policy $\pi_{\mathrm{MaxEnt}}^{*}$ has a terminating state distribution biased towards states with more complete trajectories leading to them. However, there are some situations where all the states have an equal number of trajectories leading to them. This is the case of the discrete factor graphs environments studied in Section 5.1 for example, where all terminating states can be accessed with exactly $n!$ trajectories since each of the $n$ variables can be assigned a value in any order. In this situation, one can apply MaxEnt RL with the original reward in Equation 2, without any correction, and still obtain a terminating state distribution equal to the Gibbs distribution, since this constant can be absorbed into the partition function. Just like the case where ${\mathcal{G}}$ is a tree, this must be considered as a special case though, and it is generally recommended to always correct the reward with $P_{B}$.
@@
In all applications of GFlowNets involving the SubTB objective [Madan et al., 2023, Malkin et al., 2022] found in the literature, the forward transition probabilities $P_{F}^{\theta}$ and the state flow function $F_{\phi}$ have always been parametrized with separate models (with possibly a shared backbone) as in Section 3.2. However, thanks to the equivalence with PCL established in Proposition 3.2, it is actually possible to parametrize both functions using a single “Q-function” thanks to the Unified PCL perspective [Nachum et al., 2017]; see also Section B.2 in the context of DB & SQL. A setting with separate networks is closer to the use of a dueling architecture as highlighted by Tiapkin et al. [2024]. In our experiments, we observed that using DB with two separate networks performs overall similarly to SQL with a single Q-network, even though the latter requires fewer parameters to learn. A complete study of the effectiveness of this strategy compared to separate networks is left as future work.
@@
While this paper focused on discrete distributions, to match the well-studied setting in the GFlowNet literature, the reward correction in Theorem 3.1 may be extended to cases where ${\mathcal{X}}$, along with the state and action spaces of the soft MDP, are continuous spaces as conjectured by Tiapkin et al. [2024]. We could establish similar connections with MaxEnt RL objectives by leveraging the extensions of GFlowNets to continuous spaces [Li et al., 2023, Lahlou et al., 2023]. Interestingly, continuous GFlowNets have strong connections with diffusion models [Zhang et al., 2022a, Sendera et al., 2024], for which RL has been shown to be an effective training method [Fan and Lee, 2023, Black et al., 2024].
7 Conclusion & Future work
In this work, we showed that many of the well established objectives from the GFlowNet literature [Bengio et al., 2023, Malkin et al., 2022, Pan et al., 2023a] happen to be equivalent to well-known algorithms in MaxEnt RL. Our work is anchored in the recent line of work drawing connections between GFlowNets and MaxEnt RL [Tiapkin et al., 2024, Mohammadpour et al., 2024, Niu et al., 2024], extending the reward correction introduced by Tiapkin et al. [2024] to be applicable at the level of complete trajectories. This generalization is significant as it allowed us to establish new connections between MaxEnt RL algorithms and GFlowNet objectives, especially in the presence of intermediate rewards in the underlying MDP.
This perspective makes it possible to integrate in a principled way all the tools from Reinforcement Learning for probabilistic inference over large-scale discrete and structured spaces. Future work should be dedicated to further investigating this intersection, borrowing best practices from the RL literature to enable more efficient inference in that setting. In particular, being able to better explore the state space in order to find modes of the Gibbs distribution [Rector-Brooks et al., 2023], for example using targeted exploration strategies developed for RL agents [Bellemare et al., 2016, Pathak et al., 2017], would be essential for large-scale applications such as causal structure learning [Deleu et al., 2022] and molecule generation [Bengio et al., 2021], going beyond the capabilities of MaxEnt RL alone in terms of exploration [Haarnoja et al., 2017].
@@ Acknowledgements
We would like to thank Valentin Thomas, Michal Valko, Pierre Ménard, Daniil Tiapkin, and Sobhan Mohammadpour for helpful discussions and comments about this paper. Doina Precup and Yoshua Bengio are CIFAR Senior Fellows. This research was enabled in part by compute resources and software provided by Mila (mila.quebec).
@@ Appendix A Reward correction
Recall from Ziebart [2010], Haarnoja et al. [2017] that the optimal policy maximizing LABEL:eq:maxent-rl-problem is
@@ Appendix B Equivalences between MaxEnt RL & GFlowNet objectives
SQL [Haarnoja et al., 2017] FL-DB [Pan et al., 2023a] Proposition B.3 App. B.4
$\pi$-SQL (Section 3.3) M-DB [Deleu et al., 2022] Proposition 3.3 App. B.3
@@ C.1 Probabilistic inference over discrete factor graphs
Buesing et al. [2020] defined an intermediate reward function corresponding to a decomposition of the energy Equation 73 as ${\mathcal{E}}({\bm{v}})=\sum_{t=0}^{d-1}{\mathcal{E}}(s_{t}\rightarrow s_{t+1})$, where ${\bm{v}}$ is the terminating state of a complete trajectory $(s_{0},s_{1},\ldots,s_{d},s_{f})$ (i.e., $s_{d}={\bm{v}}$), and with the partial energies defined as
${\mathcal{E}}(s_{t}\rightarrow s_{t+1})=-\sum_{m=1}^{M}\psi_{m}({\bm{v}}_{[m]})\mathbbm{1}(i\in[m]\ \mathrm{\&}\ {\bm{v}}_{[m]}\subseteq s_{t+1}),$ (74)
if the transition $s_{t}\rightarrow s_{t+1}$ corresponds to assigning the value of a particular variable $V_{i}$. In other words, the partial energy corresponds to computing all the factors as soon as all the necessary information is available (i.e., all the values of the input variables of the factors have been assigned, and $V_{i}$ that was just assigned a value is one of the input variables of the factors).
@@ C.2 Structure learning of Bayesian Networks
A state of the soft MDP corresponds to a DAG $G$ over $d$ nodes, and the initial state is the empty graph over $d$ nodes. An action consists in adding a directed edge between two nodes, such that it is not already present in the graph, and it doesn’t introduce a cycle, garanteeing that all the states of the MDP are valid acyclic graphs; there is a special action indicating whether we want to terminate and transition to $s_{f}$. Since all the states are valid DAGs, this means that all the states of the MDP are terminating, and we can use $r(G_{t},G_{t+1})={\mathcal{E}}(G_{t})-{\mathcal{E}}(G_{t+1})$ as the intermediate reward, with the appropriate energy function ${\mathcal{E}}(G)=\log P(G_{0},{\mathcal{D}})-\log P(G,{\mathcal{D}})$. Deleu et al. [2022] showed that the difference in energies can be efficiently computed using the delta-score [Friedman and Koller, 2003].
@@ C.3 Phylogenetic tree generation
To decompose the energy function into ${\mathcal{E}}(T)=\sum_{t=0}^{d-2}{\mathcal{E}}(s_{t}\rightarrow s_{t+1})$, where $T$ is the terminating state of a complete trajectory $(s_{0},s_{1},\ldots,s_{d-1},s_{f})$, we can use the observation made by Zhou et al. [2024] that the total number of mutations $M(T\mid{\bm{Y}})$ can be decomposed as the sum of (1) the number of mutations are the root of the tree, and (2) the total number of mutations in the left and right subtrees. Therefore, we can use the number of mutations at the new root of the tree constructed during the transition $s_{t}\rightarrow s_{t+1}$ as the intermediate energy ${\mathcal{E}}(s_{t}\rightarrow s_{t+1})$ (appropriately rescaled by $C$), which can be computed using the Fitch algorithm [Zhou et al., 2024].